OpenStudy (anonymous):

(3^-1 + 3^-2) ^-1 I do not understand how to work this problem.

OpenStudy (anonymous):

learning about laws on indices, powers. something ^-1 inverses it, that is to put it under 1, so 3^-1 would be 1/3 = 0.33333 something ^-2 inverses it, that is to put it under 1, and also the squares it, so 3^-2 = 1/3^2 ==> 1/9. theyre in the bracket so do them first, 1/3 + 1/9 so 3/9 + 1/9 = 4/9 so (4/9)^-1, inverses it, put one over it, but if its already a fraction, just flip the fraction so 9/4 is the answer, so 2.25.